OpenStudy (anonymous):

factor completely 25x^2+240x-100

OpenStudy (cwrw238):

the first step is to divide by the GCF of the coefficients 25 isn't a factor of 240 the GCF is 5 we have 5(5x^2 + 48x - 20)

OpenStudy (cwrw238):

now you can try to factor the expression in the parentheses)

OpenStudy (anonymous):

do u multiply 5 n -20

OpenStudy (anonymous):

is the answer 5(x-2)(x+50)

OpenStudy (cwrw238):

yes - that's a start

OpenStudy (cwrw238):

right

OpenStudy (anonymous):

thank u

OpenStudy (cwrw238):

yw

OpenStudy (anonymous):

can u help again

OpenStudy (cwrw238):

no that's wrong - (whats wrong with me today!!!!)

OpenStudy (anonymous):

5x^4-44x^2+63

OpenStudy (compassionate):

Meow. What's going on here? ~

OpenStudy (anonymous):

wats wrong with it

OpenStudy (anonymous):

tryin to find an answer to a problem

OpenStudy (cwrw238):

correct is 5(5x - 2)(x + 10)

OpenStudy (compassionate):

5x^2 + 48x - 20 can be factored further. That's what's wrong with it @cwrw238

OpenStudy (cwrw238):

yes

OpenStudy (anonymous):

can 5(x^2-13x+42 be factored more

OpenStudy (compassionate):

Yes. He just factored it for you.

OpenStudy (anonymous):

how do u do that

OpenStudy (compassionate):

Ever heard of FOIL or Binomials?

OpenStudy (anonymous):

yea

OpenStudy (cwrw238):

you can factor x^2 - 13x + 42 - you need 2 numbers a and b where ab=42 and a + b = -13

OpenStudy (cwrw238):

the factors will be of the form (x + a)( x + b)

OpenStudy (cwrw238):

what factors of 42 will ad up to -13?

OpenStudy (anonymous):

5(x-6)(x-7) is that the factored answer

OpenStudy (cwrw238):

yes

OpenStudy (anonymous):

after u do factor n sum on 5(5x^2+48-20) wat do u do

OpenStudy (cwrw238):

I don't understand that question - and do u mean -48x?

OpenStudy (cwrw238):

sorry + 48x

OpenStudy (anonymous):

just explain (25x^2+240x-20) to me

OpenStudy (anonymous):

i mean -100

OpenStudy (anonymous):

(25x^2+240x-100)

OpenStudy (cwrw238):

well first you take out the 5 - do you see that 5 is the greatest common factor of the 3 numbers?

OpenStudy (anonymous):

ya wat after that

OpenStudy (cwrw238):

now you factor 5x^2 + 48x -20 first multiply 5 * -20 = -100 with +48 in mind we see that +50 and -2 are the factors that will fit so we can write 5x^2 +50x - 2x -20 now factor by grouping: 5x(x + 10) -2(x + 10 = (5x - 2)(x + 10)

OpenStudy (cwrw238):

ok?

OpenStudy (anonymous):

thanks!

OpenStudy (cwrw238):

ok

OpenStudy (anonymous):

with (5x^4-44x^2+63) do u do it the same way